# Regional Weather Variable Predictions by Machine Learning with Near-Surface Observational and Atmospheric Numerical Data*

### v1

@@ II-C Transformer-based Models for Long-term Weather Forecasting
As the third line of work, various transformer-based solutions have been developed for long time-series predictions, including weather forecasting but in coarse resolutions. Specifically, Autoformer [24] performs weather forecasting in the daily resolution, whereas FEDformer [25] evaluates the predictions of weather time-series data for the hourly resolution. Likewise, the Corrformer model [26] forecasts weather conditions over a large number of stations in coarse granularity temporally, FourCastNet [27] considers weather predictions at the temporal resolution of 6 hours, and the iTransformer model [41] makes long-term weather forecasting in the range of 96 hours to 720 hours at the hourly resolution. Lately, general circulation models for weather and climate by combining atmospheric physics with machine learning have aimed at the daily (or longer) resolution coarsely over medium range (1-14 days) time horizons [42], which are also the target of the recently published global weather forecasting benchmark [43]. All aforementioned solutions are not meant to predict weather parameters accurately in fine temporal granularity (in minutes) over flexible time horizons and lead times, hence calling for accurate weather forecasting with fine-grained temporal resolutions.
@@ II-D Station Forecasting
The fourth line of work explores weather forecasting with several thousand weather stations as input to a forecasting model. Specifically, the Weather-5k dataset [44] presents a reliable and robust collection of over 5000 individual weather stations for evaluating multiple crucial weather elements. In addition to datasets like Weather-5k, recent advances focus on improving the accuracy of weather predictions at specific station locations through advanced downscaling techniques. A notable contribution in this area is the introduction of station-scale downscaling [45], which directly derives accurate meteorological states at station locations from coarse-resolution meteorological fields. Moreover, an innovative hybrid approach, DeepPhysiNet [46], integrates physical laws into deep learning models to enhance weather prediction accuracy. The aforementioned pursuits, however, lack predictions at fine-grained temporal resolutions, made possible only by employing near-surface parameter readings gathered with high temporal resolutions via weather stations plus geo-aligned numerical WRF-HRRR outputs.
@@ 98 Low cloud cover (%)
To compare the MiMa modelets, two other models are included in our evaluation: (1) the Micro model, (2) SARIMA, which is an autoregressive model supporting the direct modeling of the seasonal component of series  [55]. The Micro model is structured as shown in Fig. 3, taking just the micro dataset (i.e., gathered near-surface) for training an LSTM with 256 hidden states. The computed atmospheric results of the WRF-HRRR model serve as the coarser prediction counterpart (in the hourly granularity). Note that other models were included for comparison with MiMa modelets in our preliminary study [1], including support vector machine (SVR)  [56], simple neural networks (SNNs) [30], and deep uncertainty quantication models (DUQ) [43]. They were demonstrated to underperform markedly [1] and are thus not included for comparison here.

### v2

@@ II-C Transformer-based Models for Long-term Weather Forecasting
As the third line of work, various transformer-based solutions have been developed for long time-series predictions, including weather forecasting but in coarse resolutions. Specifically, Autoformer [24] performs weather forecasting in the daily resolution, whereas FEDformer [25] evaluates the predictions of weather time-series data for the hourly resolution. Likewise, the Corrformer model [26] forecasts weather conditions over a large number of stations in coarse granularity temporally, FourCastNet [27] considers weather predictions at the temporal resolution of 6 hours, and the iTransformer model [41] makes long-term weather forecasting in the range of 96 hours to 720 hours at the hourly resolution. Lately, general circulation models for weather and climate by combining atmospheric physics with machine learning have aimed at the daily (or longer) resolution coarsely over medium range (1-14 days) time horizons [42], which are also the target of the recently published global weather forecasting benchmark [43].
@@ II-D Station Forecasting
The fourth line of work includes weather forecasting via (1) data from a large number of weather stations [44], (2) station downscaling [45], (3) a physical-ML hybrid model [46], (4) data from dense and sparse sensors [47], and (5) GNN-based method [48]. Specifically, multiple critical weather variables are forecast using a comprehensive collection of over 5,000 weather stations, called the Weather-5k dataset [44]. Leveraging the dataset like Weather-5k, recent pursuits improve weather predictions at specific station locations through advanced downscaling techniques. For instance, station-scale downscaling [45] accurately derives meteorological conditions at station locations from coarse-resolution meteorological fields. Meanwhile, the hybrid DeepPhysiNet approach [46] integrates physical laws into deep learning models to enhance prediction accuracy. Another study [47] utilizes both dense and sparse sensor data to make predictions for lead times up to 24 hours, with its focus on extending the lead time from 12 to 24 hours. However, the study relies solely on ground-based sensor data without valuable computational numerical data, such as WRF-HRRR, which would enhance accuracy. A GNN-based method [48] has also been proposed for downscaling global grids to off-grid locations of interest. However, like the previous study, it does not leverage computational numerical data, such as WRF-HRRR, to improve predictions for these off-grid locations. While all aforementioned solutions help to advance weather prediction, they are not meant to predict weather parameters accurately in fine temporal granularity (in minutes) over flexible time horizons and lead times, hence calling for accurate weather forecasting with fine-grained temporal resolutions.
@@ 98 Low cloud cover (%)
To compare the MiMa modelets, five other models are included in our evaluation: (1) the Micro model, (2) SARIMA, which is an autoregressive model supporting the direct modeling of the seasonal component of series  [57], (3) SNN, which is a simple neural network [30] that takes the Micro data to make predictions, (4) SVR, which is a regression model based on support vector machines [58], with the Micro data as input features, and (5) DUQ (Deep Uncertainty Quantification), which integrates deep learning techniques with one GRU layer of 128 hidden nodes [43] to quantify uncertainties in weather predictions, enhancing the reliability of weather forecasts. The Micro model is structured as shown in Fig. 3, taking just the micro dataset (i.e., gathered near-surface) for training an LSTM with 256 hidden states. The computed atmospheric results of the WRF-HRRR model serve as the coarser prediction counterpart (in the hourly granularity). While these comparative models are evaluated alongside MiMa modelets, they consistently underperform as demonstrated in both our preliminary study [1] and the results presented in Tables VI and XI below.
@@ Acknowledgement
SNN 2.63 (9.61%) 1.72 (6.26%) 8.64 (11.83%) 5.03 (6.88%) 0.65 (6.86%) 0.42 (4.41%) 220.93 (39.45%) 203.31 (36.31%)
SVR 3.41 (12.43%) 2.50 (9.12%) 9.99 (13.67%) 6.60 (9.04%) 0.80 (8.37%) 0.62 (6.55%) 48.88 (8.73%) 33.05 (5.90%)
DUQ 2.19 (12.18%) 1.65 (9.18%) 7.29 (12.85%) 4.27 (7.53%) 1.21 (12.48%) 0.86 (8.89%) 102.35 (1052.98%) 100.45 (1033.50%)
Station Model TEMP HUMI WSPD PRES
RMSE MAE RMSE MAE RMSE MAE RMSE MAE